OpenStudy (anonymous):

I need help with algebra 2! Solve the equation. Check for extraneous solutions. 2|7-7x|=2x+4

OpenStudy (anonymous):

Solve each equation. Check for extraneous solutions. Equation: |5x-1|+7=3x Here's what I did: |5x-1|+7=3x -7 -7 |5x+1|=3x-7 5x-1=3x-7 OR 5x-1=-(3x-7) +1 +1 5x=3x-6 -3x -3x 2x=-6 2 2 x=-3 OR 5x-1=-(3x-7) 5x-1=-3x+7 8x = 8 x = 1 So, x=-3 or x=1. CHECK: |5x-1|+7=3x

OpenStudy (anonymous):

|5(-3)-1|+7=+3(-3) |-15-1|+7=-9 |-16|+7=-9 16+7=-9 23=9 But 23 doesn't equal 9. At first I thought it was just an exraneous solution. But then I checked the second number: |5x-1|+7=3x |5(1)-1|+7=3(1) |5-1|+7=3 |4|+7=3 4+7=3 11=3 But again, 11 doesn't equal 3. Your work is correct. Both "solution" are extraneous. --- If you graph the left side and the right side of the equation separately, you will see that they do not intersect, i.e. they are never equal.

OpenStudy (anonymous):

Its not graphing

OpenStudy (anonymous):

The answers are a. x=3/2 x=5/6 b. x=3/2 x=5/8 c. x=5/8 d. x=3/2

OpenStudy (anonymous):

A. x=3/2 x=5/6 is the correct 1

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

no prob gimme a medal